# Gravitational waves from cosmic strings in LISA: reconstruction pipeline and physics interpretation

### v1

@@ 4.4 Different injection and recovery templates
Our tests up to this point have been done by injecting a signal using some model with fixed parameters, then fitting the signal using that same model. This is a good test of the reconstruction capabilities, but may not be entirely accurate to the real situation of fitting an unknown cosmic string background with a template. Model uncertainties and approximations made in creating the SGWBs used here mean that the true string signal will likely be slightly different from our templates — the question of the effect of gravitational backreaction on the SGWB has already been mentioned, as one example.
To get a first-order sense of the reconstruction performance in such a scenario, we take advantage of one of its features: the template used to inject the signal and the template used to recover (or fit) the signal do not have to be the same. For our example, we will inject a Model II signal at a fixed tension and fit it using Model I.
Let us begin with a mild test of the typical approximation of Model II by Model I where one takes $\alpha=0.1$, $q=4/3$, and $\mathcal{F}=0.1$. In this case, the amplitude of the radiation era plateau predicted by both models is asymptotically identical. However, since the power spectrum is modelled in different ways in these models, the shape of the peak of the SGWB may be slightly different. Moreover, Model I, unlike Model II, includes the impact of the evolution of the effective number of DoF in the dynamics of the network and in loop production, which results in a smoothing of these features. As a result, the matching between these two models for this set of parameters, although quite good, is not perfect. Fixing those three parameters — allowing only $G\mu$ to vary — and including foregrounds, we perform 30 recoveries of a Model II signal with $\log_{10}(G\mu)=-14.0$ by Model I. The average and standard deviation of the recovered logarithmic tension are $-13.893$ and $1.853\times 10^{-2}$, respectively. For reference, the fitting of Model II by Model II (as shown in figures 7, 8) at $\log_{10}(G\mu)=-14.0$ gave a mean logarithmic tension of $-13.998$ with a standard deviation of $2.498\times 10^{-2}$.
While the relative difference in the actual tensions is small (about $+28$%), and may not be significant compared to theoretical errors in the predicted energy scale and couplings in a BSM theory containing strings, the recovered logarithmic tension here is $\sim 5.8$ standard deviations away from the injected logarithmic tension. Put another way, for this single-parameter fit at moderate tensions, any reasonable statistical test would reject the “true” tension in this case in favour of a slightly heavier string.
This discrepancy is not a sign of any serious incompatibility between Models I and II, but rather should be understood as arising from the slight mismatch between the SGWB, especially around the degree-of-freedom steps, as discussed above. A secondary effect comes from the canonical values of $\alpha$, $q=$, and $\mathcal{F}$ used when comparing the models, which may be slightly different from the “ideal” values from SGWBinner’s perspective. For example, ref. Blanco-Pillado:2011egf gives the peak of loop size distribution at creation as (in the language of this paper) $\alpha=0.09$ in the radiation era. Combined with a similarly-sized difference in $\mathcal{F}$, this can explain the discrepancy. [^22] The choice of $q=4/3$ is theoretically well-motivated by saying that the GW spectrum is dominated by cusp emission, but a slight difference in the best value to use to fit Model II could be explained as accounting for the low-mode difference between the Model II $P_{j}$ and a pure power law $P_{j}$.
We can also investigate how well Model I fits Model II with no parameters fixed; that is, can SGWBinner arrive at the correct tension and the typical approximation values starting from a completely naïve search? The result, unsurprisingly, is regime-dependent. Again using an injected $\log_{10}(G\mu)=-14.0$ and including foregrounds, we find the means and standard deviations of the Model I template’s parameters as shown in the first two rows of table 2.
@@
Injected $\log_{10}(G\mu)=-14.0$ Mean $-13.6$ $-1.46$ $1.51$ $-0.851$
Std. dev. $0.384$ $0.338$ $0.0998$ $0.151$
Injected $\log_{10}(G\mu)=-11.0$ Mean $-12.6$ $-0.832$ $1.23$ $-0.404$
Std. dev. $0.365$ $0.212$ $0.0228$ $0.0974$
@@
Now, the recovered tension falls just above one standard deviation from the actual tension, and the log-fuzziness just below. The log-$\alpha$ and spectral index fall within two standard deviations. For intermediate tensions, the recovery of a near-but-not-exact signal by a Model I template can be given moderate credence, although further testing and optimization of the code is necessary to obtain robust and reliable results.
@@
Larger tensions run into issues of degeneracy, whereas at lower tensions the foregrounds cause a degradation of fitting quality regardless of what template recovers what signal. By way of example, we repeat our process for an injected $\log_{10}(G\mu)=-11.0$, as shown in the last two rows of Table 2. Here, much of the mismatch between the expected outcomes and the recovered values can be explained by the degeneracy between the Model I parameters at large tension values. In the case studied, the log-fuzziness parameter is notably higher than the expected $-1.0$, raising the height of the SGWB, but the tension is notably lesser, lowering the height of the SGWB. Because an SGWB at this tension is high above the LISA noise curve and foregrounds, SGWBinner reports very low uncertainties on these fits, making the difference between the recovered and injected parameters (in a number-of-standard-deviations sense) much worse than the lower-tension case we first studied.
These results can be improved with a more thorough understanding of the degeneracies between Model I parameters for an SGWB measured in the LISA band. While these degeneracies have already been pointed out for large tensions, intermediate (and low) tensions are not exempt; in particular, a good starting point for better understanding is looking closely at the parameter-space region around $\left\{\log_{10}(\alpha),\,q,\,\log_{10}(\mathcal{F})\right\}=\left\{-1,4/3,-1\right\}$, as we would expect a “standard” cosmic string signal to appear in this vicinity.
@@ 5 Science interpretation
Finding a stochastic background of GWs consistent with one of the templates described in this paper would have a huge impact on the study of physics BSM.
@@
On the other hand, current numerical studies of Nambu-Goto cosmic string networks seem to indicate that the scaling distribution of non-intersecting loops peaks at around $\alpha\sim 0.1$ Blanco-Pillado:2013qja . A substantial departure from this regime could be, in principle, identified using a Model I reconstruction. Finding a large departure from the preferred value of $\alpha$ would likely point to a more complex cosmic string model than the one discussed here. Numerical simulations of strings with reduced intercommutation probability Avgoustidis:2005nv , for instance, suggest that the production of small loops may be favoured in cosmic superstring networks. Such a departure from the expected values of $\alpha$ could also hint at the possibility of a much richer string micro-physics than that of the strings studied in this paper. Models with such new internal structure have been suggested in the literature  Witten:1984eb ; Nielsen:1987fy ; Carter:1990nb ; Vilenkin:1990mz , but the details of loop production and their implications in terms of their GW signatures are much less understood at this point.
The biggest theoretical uncertainty at the moment continues to be the form of the power spectrum from individual strings, the so-called $P_{j}$ spectrum. Future investigations using an accurate model of gravitational backreaction would undoubtedly improve the situation before the arrival of the LISA data. On the other hand, these models also assume a somewhat minimalistic version of the string microphysics, therefore, obtaining a good handle on the slope of this spectrum (our $q$ parameter) could be helpful in identifying possible deviations from these simple models. It is therefore interesting to point out the ability of our reconstructions to limit the possible values of the exponent of a power-law fit to this spectrum.
@@ Appendix A Complete template for Model I
In this appendix, we provide a detailed derivation of the analytical template for Model I, including the contributions of all relevant modes of emission. To do so, we resort to the full analytical approximation to the contribution of the $j$-th mode of emission to the SGWB derived in ref. Sousa:2020sxs , which may be re-written as  [^23] Note that here, as explained in section 2.2, we re-express the results of ref. Sousa:2020sxs in terms of $\alpha$ instead of using the natural parameter of the model $\alpha_{L}$. Moreover, the constants in eq. (4) have been absorbed into the definition of $C_{r}$, $C_{m}$ and $C_{rm}$, as in ref. Sousa:2020sxs .
@@
where $\mathrm{\Omega}_{{\rm gw}}^{r\,j}$ represents the contribution of the loops that are created and decay in the radiation era, $\mathrm{\Omega}_{{\rm gw}}^{rm\,j}$ that of radiation-era loops that survive into the matter era and $\mathrm{\Omega}_{{\rm gw}}^{m\,j}$ corresponds to the contribution of the loops that are created in the matter era. Here, we defined
$\begin{gathered}C_{r}=\frac{128}{9}\pi\mathcal{C}_{r}\Omega_{r}\left(1+\epsilon_{r}\right)^{3/2}\frac{G\mu}{\epsilon_{r}},\quad C_{rm}=32\sqrt{3}\pi\mathcal{C}_{r}\left(1+\epsilon_{r}\right)^{3/2}\frac{H_{0}\left(\Omega_{m}\Omega_{r}\right)^{3/4}}{\epsilon_{r}\Gamma f}\,,\\ C_{m}=162\pi\mathcal{C}_{m}\frac{(1+\epsilon_{m})}{G\mu\epsilon_{m}}\left(\frac{H_{0}\Omega_{m}}{\Gamma f}\right)^{2}\,,\quad\mathcal{C}_{i}=\frac{\tilde{c}}{\sqrt{2}}\mathcal{F}\frac{v_{i}}{\xi_{i}^{3}}\,,\quad D_{i}=\frac{2H_{0}\Omega_{i}^{1/2}}{\nu_{i}\Gamma G\mu}\,,\\ {\tilde{A}_{r}}=\frac{D_{r}}{f}\frac{\Omega_{m}}{\Omega_{r}}\,,\quad{\tilde{A}_{rm}}=\frac{D_{m}}{f}\sqrt{\frac{\Omega_{m}}{\Omega_{r}}}\,,\quad{\tilde{A}_{m}}=\frac{D_{m}}{f}\,,\quad{\tilde{A}}_{0}=\frac{2\sqrt{2}}{f\Gamma}\sqrt{\frac{H_{0}\Omega_{r}^{1/2}}{t_{pl}\left(1+\epsilon_{r}\right)}}\,,\\ \epsilon_{r}=\frac{\alpha}{\Gamma G\mu}\,,\quad\epsilon_{m}=\epsilon_{r}\frac{\xi_{m}}{\xi_{r}}\,,\quad\gamma_{m}=1+\epsilon_{m}^{-1}\,,\quad\beta_{m}=\frac{1}{\epsilon_{m}}\left(1+\frac{1}{\gamma_{m}}\right)\,.\end{gathered}$ (36)
Labels $i=r,m$ are used to refer to the values of the corresponding variables in the radiation and matter eras, respectively. In the radiation era, for standard cosmic strings, we have $\nu_{r}=1/2$, $\xi_{r}=0.271$, $v_{r}=0.662$ and $\mathcal{C}_{r}=5.4\mathcal{F}$. While, in the matter era, $\nu_{m}=2/3$, $\xi_{m}=0.625$, $v_{m}=0.583$ and $\mathcal{C}_{m}=0.39\mathcal{F}$. Moreover, $\tilde{c}=0.23$ Martins:2000cs in both radiation and matter eras.
As explained in section 2.2, to fully characterize the SGWB generated by cosmic strings, we need to sum the contributions of all relevant harmonic modes, which implies computing the summation
(where $a=r$, $rm$ or $m$ labels the different contributions to the SGWB) up to $N(f)=f/f_{min,r}$ for $\mathrm{\Omega}_{\rm gw}^{r}$ and $\mathrm{\Omega}_{\rm gw}^{rm}$ and up to $N(f)=f/f_{min,m}$ in the case of $\mathrm{\Omega}_{\rm gw}^{m}$ (where $f_{min,r}=D_{r}\Omega_{m}/(\epsilon_{r}\Omega_{r})$ and $f_{min,m}=D_{m}/\epsilon_{m}$). In order to do so, we resort to the Euler–Maclaurin summation formula, which allows us to obtain fast and accurate results without performing the summation explicitly. According to this approximation knopp1990theory we should have that:
@@
The integration of the contributions in eqs. (33)-(35) can be carried out analytically by using the following relation
[…]
where ${}_{2}F_{1}(\dots)$ is the Gauss hypergeometric function. To compute the necessary derivatives, it is useful to define
@@
Hence, including terms up to the third derivative, we obtain the following approximation:
@@
As to the term in $\mathrm{\Omega}_{\rm gw}^{m}$ involving a logarithm, we can use (39) and integration by parts:
@@
Hence, the summation involving this term can be approximated by:
@@
Using the notation introduced above and the Euler-Maclaurin approximation (38), we may express the radiation, radiation-matter and matter contributions to the SGWB — including the contribution of all relevant harmonic modes — respectively as
@@
Equation (47) does not include the impact of the change of the effective number of massless DoF as the universe expands. We can take this effect into consideration by breaking the integration in ref. Sousa:2020sxs into smaller intervals in which the measure of the number of massless DoF, $\Delta_{r}(a)$ (defined in eq. (7)), remains roughly constant:
$\begin{gathered}\mathrm{\Omega}_{\rm gw}^{r}(f,\alpha,G\mu)=\frac{128}{9}\pi\left(\frac{G\mu}{\Gamma}\right)^{1/2}\Omega_{r}C_{r}(\alpha)\Bigg{[}\Delta_{r\,1}\left(\frac{1}{B_{1}(a)+1}\right)^{3/2}\Bigg{|}_{a=a_{0}^{*}}^{a=a^{*}_{1}}\\ +\Delta_{r\,2}\left(\frac{1}{B_{2}(a)+1}\right)^{3/2}\Bigg{|}_{a=a^{*}_{1}}^{a=a^{*}_{2}}+\dots+\Delta_{r\,N_{\text{dof}+1}}\left(\frac{1}{B_{N_{\text{dof}+1}}(a)+1}\right)^{3/2}\Bigg{|}_{a=a^{*}_{N_{\text{dof}}}}^{a=a_{\text{eq}}}\Bigg{]},\end{gathered}$ (49)
where we defined $B_{n}(a)\equiv D_{r}\Delta^{1/2}_{r\,n}(a_{0}/a)$ and $C_{r}(\alpha)\equiv\mathcal{C}_{r}\left(\alpha+\Gamma G\mu\right)^{3/2}/\alpha$. Here we have approximated the changes in the effective number of relativistic DoF as a piecewise constant function with $N_{\text{dof}}$ steps occurring for values of scale factor $a^{*}_{1},\cdots,a^{*}_{N_{\text{dof}}}$. $\Delta_{r\,n}$ is the (constant) value of $\Delta_{r}(a)$ between $a^{*}_{n-1}$ and $a^{*}_{n}$, while $a^{*}_{N_{\text{dof}}+1}=a_{\text{eq}}$ and the start of integration corresponds to $a^{*}_{0}=a_{0}(2H_{0}\Omega_{r}^{1/2}(\epsilon_{r}+1)t_{\text{pl}})^{1/2}/(G\mu)$ (corresponding to the start of significant GW emission from cosmic string loops). Using the Euler–Maclaurin summation formula, we obtain the following expression for the contribution of loops that decay during the radiation era:

### v2

@@ 4.4 Different injection and recovery templates
One can also ask if the LISA SGWB search needs to include both Model I and Model II in the template library. Model I with fixed parameters $\alpha=0.1$, $\mathcal{F}=0.1$, and $q=4/3$ is typically taken as a good approximation of Model II. However, while the agreement at the high-frequency plateau is exact, there are important, frequency-dependent differences between these models. Since the power spectrum is modelled in different ways in these models, the shape of the peak of the SGWB can be slightly different. Moreover, Model I, unlike Model II, includes the impact of the evolution of the effective number of DoF in the dynamics of the network and in loop production, which results in a smoothing of these features. As a result, the matching between these two models for this set of parameters, although quite good, is not perfect. Nevertheless, such a mismatch is not necessarily relevant for LISA. If the mismatch between the two templates was largely within the reconstruction error bars in the whole cosmic string parameter space, one could disregard one of the templates in the frequency shape reconstruction required to isolate the SGWB signal in the global fit.
@@
Moreover, our tests up to this point have been done by injecting a signal using some model with fixed parameters, and then fitting the signal using that same model. This is a good test of the reconstruction capabilities, but may not be entirely accurate to the real situation of fitting an unknown cosmic string background with a template. Theoretical uncertainties on the cosmic string modelling might indeed introduce systematics, meaning that the true string signal might be slightly different from our templates — the question of the effect of gravitational backreaction on the SGWB has already been mentioned, as one example.
To investigate both of these questions, we take advantage of one of SGWBinner’s features: the template used to inject the signal and the template used to recover (or fit) the signal does not have to be the same. For our example, we will inject a Model II signal at a fixed tension and fit it twice, once using Model II and once using Model I. We will not fix any of the parameters of Model I, as we wish to see if it naturally selects the typical values of $\alpha$, $\mathcal{F}$, and $q$ mentioned above. We will then investigate both the recovered parameters of the fit using Model I, and compare the evidences for the fit using Model II versus using Model I.
We inject Model II in the presence of foregrounds with two different tensions: $\log_{10}(G\mu)=-11.0$ and $-14.0$. The first is done to test the reconstruction in a regime where most of the LISA band sees the high-frequency plateau, where the agreement should be exact; the second is done to see how the reconstruction is affected for the tensions where the foreground’s overlap with the string GWB becomes significant (see Fig. 8 and nearby discussion). Thirty injections are performed for each tension, and the result of the reconstructions are summarized in table 2. Larger tensions run into issues of degeneracy, whereas at lower tensions the foregrounds cause a degradation of fitting quality regardless of what template recovers what signal.
@@
Injected $\log_{10}(G\mu)=-11.0$ Mean $-12.1$ $-1.05$ $1.33$ $-0.484$
Std. dev. $0.0842$ $0.608$ $0.00130$ $0.297$
Injected $\log_{10}(G\mu)=-14.0$ Mean $-14.0$ $-1.12$ $1.53$ $-0.896$
Std. dev. $0.538$ $0.563$ $0.149$ $0.318$
@@
The recovery for injected $\log_{10}(G\mu)=-11.0$ is shown in the first two rows of Table 2. The mean recovered tension is many standard deviations below the injected tension; $\log_{10}(\alpha)$ and $q$ are within one standard deviation of the typical values used when approximating a Model II SGWB using Model I, but the log-fuzziness is just over two standard deviations too high. This indicates that it would not be correct to perform searches using only Model I templates, as a Model II-like signal would be misidentified as coming from a network with different parameters and thus different physics.
@@
Here, much of the mismatch between the expected outcomes and the recovered values can be explained by the degeneracy between the Model I parameters at large tension values. In the case studied, the log-fuzziness parameter is higher than the expected $-1.0$, raising the height of the SGWB, but the tension is lesser, lowering the height of the SGWB. Because an SGWB at this tension is high above the LISA noise curve and foregrounds, SGWBinner reports very low uncertainties on these fits, making the difference between the recovered and injected parameters (in a number-of-standard-deviations sense) much worse than the lower-tension case we first studied.
For $\log_{10}(G\mu)=-14.0$, the recovered tension, log-fuzziness, and log-loop size all fall within one standard deviation of their expected or typical values; the power spectral index about one and a half standard deviations high. For intermediate tensions, the recovery of a near-but-not-exact signal by a Model I template can be given moderate credence. However, further testing and optimization of the code is necessary to obtain robust and reliable results—note, for example, the significant increase in the standard deviation of the fit $\log_{10}(G\mu)$ and $q$ values between log-tensions of $-11.0$ and $-14.0$.
[…]
Finally, we may quantify which model is a better fit for any given reconstruction using the evidences reported by polychord, used by SGWBinner to find best-fit parameters in each run. For each of the 30 runs mentioned above, we found the log-Bayes factor with the convention that $\log_{10}(BF)>0$ indicates support for fitting the signal with Model II. For the $\log_{10}(G\mu)=-11.0$ injections, we found the mean of $\log_{10}(BF)$ to be $7.61$, with a range of $5.32$ to $10.3$. In all cases, this is decisive evidence for reconstructing the signal with Model II. For the $\log_{10}(G\mu)=-14.0$ injections, we found the mean of $\log_{10}(BF)$ to be $3.74$, with a range of $0.113$ to $5.99$; while Model II was still preferred in every scenario, in one realization the evidence is not worth mentioning, and only 27 of the 30 realizations contain decisive evidence for Model II.
While we have not attempted a reconstruction of an injected Model I signal using both Models I and II, we anticipate similar results, if not stronger: the additional parameters in Model I permit a greater range of SGWB morphologies, and a Model II template should never match, e.g., a Model I signal with $\alpha=10^{-3}$ (except, perhaps, in the very-high-tension plateau case, as discussed above). This underscores the necessity of having both Model I and Model II in the SGWBinner template bank.
4.5 Models with non-gravitational radiation
The preceding sections looked at string models where the only channel for string decay is into GWs. Outside of these minimally-coupled strings, several models exist which have other, sometimes dominant, modes of energy loss.
Consider the work of refs. Matsunami:2019fss ; Auclair:2019jip ; Hindmarsh:2021mnl ; Hindmarsh:2022awe , which suggests modeling an Abelian–Higgs (AH) network of strings by multiplying a Nambu–Goto (NG) SGWB by a parameter $f_{\text{NG}}$, representing the non-gravitational radiation:
Interestingly, the uncertainty in the case of $\log_{10}(G\mu)=-12.0$, $f_{\text{NG}}=0.1$ is less than the other points with $f_{\text{NG}}=0.1$, but its absolute error is greater when compared to the points adjacent to it. The reconstructions of this point prefer heavier strings with lower Nambu–Goto fraction than the injected values. While it is not immediately clear why this is the case, the answer may lie in which parts of the string SGWB are in the LISA window. Much of the structure in the string SGWB is in the radiation-to-matter-transition ‘bump’ and in the changing-degree-of-freedom ‘steps’. If the portion of the SGWB in the LISA window lacks much of this structure (as one might get by simply lowering a high-tension SGWB), then a slightly heavier string which emits less of its energy into GWs might be able to provide a better fit by SGWBinner’s metrics.
@@ 5 Science interpretation
Finding a stochastic background of GWs consistent with one of the templates described in this paper would have a huge impact on the study of physics beyond the standard model.
@@
On the other hand, numerical studies of Nambu-Goto cosmic string networks that we used in this paper (Model II) seem to indicate that the scaling distribution of non-intersecting loops peaks at around $\alpha\sim 0.1$ Blanco-Pillado:2013qja . A substantial departure from this regime could be, in principle, identified using a Model I reconstruction. Finding a large departure from the preferred value of $\alpha$ would likely point to a more complex cosmic string model than the one discussed here. Numerical simulations of strings with reduced intercommutation probability Avgoustidis:2005nv , for example, suggest that the production of small loops may be favored in cosmic superstring networks. Such a departure from the expected values of $\alpha$ could also hint at the possibility of a much richer string micro-physics than that of the strings studied in this paper. Models with such new internal structure have been suggested in the literature  Witten:1984eb ; Nielsen:1987fy ; Carter:1990nb ; Vilenkin:1990mz , but the details of loop production and their implications in terms of their GW signatures are much less understood at this point. Alternatively, interpretations of such a deviation could potentially result from the possibility known in the literature as Model III Ringeval:2005kr ; Lorenz:2010sm ; Auclair:2019zoz ; Auclair:2020oww .
Observation of the spectrum would also help to settle the debate on possible emission other than GWs from the cosmic string network Vachaspati:1984yi ; Vincent:1997cx ; Moore:1998gp ; Olum:1998ag ; Olum:1999sg ; Moore:2001px ; Hindmarsh:2008dw ; Hindmarsh:2017qff ; Matsunami:2019fss ; Hindmarsh:2021mnl ; Blanco-Pillado:2022rad ; Hindmarsh:2022awe ; Blanco-Pillado:2023sap . In Sec. 4.5 we find that a mild reduction of the GW emission would still allow for the observation of the spectrum although the degeneracy of the emission efficiency with $G\mu$ makes extraction of the value more difficult in much of the parameter space.
Within our Models I and II, there are still theoretical uncertainties. The biggest oneat the moment continues to be the form of the power spectrum from individual strings, the so-called $P_{j}$ spectrum. Future investigations using an accurate model of gravitational backreaction would undoubtedly improve the situation before the arrival of the LISA data. On the other hand, these models also assume a somewhat minimalistic version of the string microphysics, therefore, obtaining a good handle on the slope of this spectrum (our $q$ parameter) could be helpful in identifying possible deviations from these simple models. It is therefore interesting to point out the ability of our reconstructions to limit the possible values of the exponent of a power-law fit to this spectrum.
@@ Appendix A Complete template for Model I
In this appendix, we provide a detailed derivation of the analytical template for Model I, including the contributions of all relevant modes of emission. Although analytical approximations to the contribution of a single harmonic mode of emission were derived in ref. Sousa:2020sxs , to fully characterize the SGWB generated by cosmic strings, we need to sum the contributions of all relevant harmonic modes, which implies computing the summation
where $a=r$, $rm$ or $m$ label respectively the contribution of the loops that are created and decay in the radiation era, that of radiation-era loops that survive into the matter era and the contribution of the loops that are created in the matter era. This summation needs to be performed, as explained in Sec. 2.2, up to $N(f)=f/f_{min,r}$ for $\mathrm{\Omega}_{\rm gw}^{r}$ and $\mathrm{\Omega}_{\rm gw}^{rm}$ and up to $N(f)=f/f_{min,m}$ in the case of $\mathrm{\Omega}_{\rm gw}^{m}$ Since performing it explicitly is time-costly, here we resort to the Euler–Maclaurin summation formula, as it allows us to obtain fast and accurate results without performing the summation explicitly. According to this approximation knopp1990theory we should have that:
[…]
We will then resort to the analytical approximation to the contribution of the $j$-th mode of emission to the SGWB derived in ref. Sousa:2020sxs to construct a novel template that includes all harmonic modes of emission using eq. (35). The essential physical ingredients of this approximation are:
  • String loop parameters: from Model I (see Table 1) we have tension $G\mu$, loop size $\alpha$, power spectral index $q$, and fuzziness $\mathcal{F}$; in addition, we require the loop rate of GW radiation $\Gamma$.
  • Parameters of the velocity one-scale (VOS) model Martins:1996jp ; Martins:2000cs : the RMS velocity of the cosmic string network, $\bar{v}$; the efficiency of the energy-loss mechanism of the cosmic string network, $\tilde{c}$; and the charateristic lengthscale of the network, $L=\xi t$.
  • Cosmological parameters: the Hubble constant $H_{0}$; the relative energy density in matter and radiation, $\Omega_{m}$ and $\Omega_{r}$; the Planck time $t_{pl}$; and the power $\nu$ in the scale factor relation $a\propto t^{\nu}$.
From these physical parameters, it is helpful to define many auxiliary terms to simplify the eventual expressions for $\Omega_{\text{gw}}$. With $f$ being the frequency, these auxiliary terms are:
$\begin{gathered}\epsilon_{r}=\frac{\alpha}{\Gamma G\mu}\,,\quad\epsilon_{m}=\epsilon_{r}\frac{\xi_{m}}{\xi_{r}}\,,\quad\gamma_{m}=1+\epsilon_{m}^{-1}\,,\quad\beta_{m}=\frac{1}{\epsilon_{m}}\left(1+\frac{1}{\gamma_{m}}\right)\,,\\ \mathcal{C}_{i}=\frac{\tilde{c}}{\sqrt{2}}\mathcal{F}\frac{v_{i}}{\xi_{i}^{3}}\,,\quad C_{m}=162\pi\mathcal{C}_{m}\frac{(1+\epsilon_{m})}{G\mu\epsilon_{m}}\left(\frac{H_{0}\Omega_{m}}{\Gamma f}\right)^{2}\,,\quad D_{i}=\frac{2H_{0}\Omega_{i}^{1/2}}{\nu_{i}\Gamma G\mu}\,,\\ C_{r}=\frac{128}{9}\pi\mathcal{C}_{r}\Omega_{r}\left(1+\epsilon_{r}\right)^{3/2}\frac{G\mu}{\epsilon_{r}},\quad C_{rm}=32\sqrt{3}\pi\mathcal{C}_{r}\left(1+\epsilon_{r}\right)^{3/2}\frac{H_{0}\left(\Omega_{m}\Omega_{r}\right)^{3/4}}{\epsilon_{r}\Gamma f}\,,\\ {\tilde{A}_{r}}=\frac{D_{r}}{f}\frac{\Omega_{m}}{\Omega_{r}}\,,\quad{\tilde{A}_{rm}}=\frac{D_{m}}{f}\sqrt{\frac{\Omega_{m}}{\Omega_{r}}}\,,\quad{\tilde{A}_{m}}=\frac{D_{m}}{f}\,,\quad{\tilde{A}}_{0}=\frac{2\sqrt{2}}{f\Gamma}\sqrt{\frac{H_{0}\Omega_{r}^{1/2}}{t_{pl}\left(1+\epsilon_{r}\right)}}\,.\end{gathered}$ (36)
Labels $i=r,m$ are used to refer to the values of the corresponding variables in the radiation and matter eras, respectively. In the radiation era, for standard cosmic strings, we have $\nu_{r}=1/2$, $\xi_{r}=0.271$, $v_{r}=0.662$ and $\mathcal{C}_{r}=5.4\mathcal{F}$, while in the matter era, $\nu_{m}=2/3$, $\xi_{m}=0.625$, $v_{m}=0.583$ and $\mathcal{C}_{m}=0.39\mathcal{F}$. Moreover, $\tilde{c}=0.23$ Martins:2000cs in both radiation and matter eras. Note also that $f_{min,r}=D_{r}\Omega_{m}/(\epsilon_{r}\Omega_{r})$ and $f_{min,m}=D_{m}/\epsilon_{m}$.
With these definitions, the results of ref. Sousa:2020sxs may be re-written as  [^21] Note that here, as explained in section 2.2, we re-express the results of ref. Sousa:2020sxs in terms of $\alpha$ instead of using the natural parameter of the model $\alpha_{L}$. Moreover, the constants in eq. (4) have been absorbed into the definition of $C_{r}$, $C_{m}$ and $C_{rm}$, as in ref. Sousa:2020sxs .
@@
where $\mathrm{\Omega}_{{\rm gw}}^{r\,j}$ represents the contribution of the loops that are created and decay in the radiation era, $\mathrm{\Omega}_{{\rm gw}}^{rm\,j}$ that of radiation-era loops that survive into the matter era and $\mathrm{\Omega}_{{\rm gw}}^{m\,j}$ corresponds to the contribution of the loops that are created in the matter era.
Note that most of the terms in eqs. (37)-(39) may be written in the form $j^{a}(1+Aj)^{-b}$ for different values of $a$, $b$ and $A$ (e.g. the first term in eq. (37) corresponds to $a=-q$, $b=3/2$ and $A=\tilde{A}_{r}$). The computation of the first term in eq. (35) then involves the computation of integrals of the form:
@@
where ${}_{2}F_{1}(\dots)$ is the Gauss hypergeometric function. Moreover, the derivatives needed to compute the last term in eq. (35) yield:
@@
Combining these results, one finds that, including terms up to the third derivative, terms of this form lead to a contribution to the full spectrum of the form:
@@
There is only one term with a distinct form in $\mathrm{\Omega}_{\rm gw}^{m}$ (the last term in eq. (39), which involves a logarithm), but in this case the integral may also be performed analytically if we use (40) and integrate by parts:
@@
This term then gives rise to a contribution of the form:
@@
We then find that the radiation, radiation-matter and matter contributions to the SGWB — including the contribution of all relevant harmonic modes — may be expressed respectively as
@@
Note however that eq. (48) does not yet include the impact of the change of the effective number of massless DoF as the universe expands. We can take this effect into consideration by breaking the integration in ref. Sousa:2020sxs into smaller intervals in which the measure of the number of massless DoF, $\Delta_{r}(a)$ (defined in eq. (7)), remains roughly constant. In this case, (37) may be rewritten as:
$\begin{gathered}\mathrm{\Omega}_{\rm gw}^{r\,1}(f,\alpha,G\mu)=\frac{128}{9}\pi\left(\frac{G\mu}{\Gamma}\right)^{1/2}\Omega_{r}C_{r}(\alpha)\Bigg{[}\Delta_{r\,1}\left(\frac{1}{B_{1}(a)+1}\right)^{3/2}\Bigg{|}_{a=a_{0}^{*}}^{a=a^{*}_{1}}\\ +\Delta_{r\,2}\left(\frac{1}{B_{2}(a)+1}\right)^{3/2}\Bigg{|}_{a=a^{*}_{1}}^{a=a^{*}_{2}}+\dots+\Delta_{r\,N_{\text{dof}+1}}\left(\frac{1}{B_{N_{\text{dof}+1}}(a)+1}\right)^{3/2}\Bigg{|}_{a=a^{*}_{N_{\text{dof}}}}^{a=a_{\text{eq}}}\Bigg{]},\end{gathered}$ (50)
where we defined $B_{n}(a)\equiv D_{r}\Delta^{1/2}_{r\,n}(a_{0}/a)$ and $C_{r}(\alpha)\equiv\mathcal{C}_{r}\left(\alpha+\Gamma G\mu\right)^{3/2}/\alpha$. Here we have approximated the changes in the effective number of relativistic DoF as a piecewise constant function with $N_{\text{dof}}$ steps occurring at values of scale factor $a^{*}_{1},\cdots,a^{*}_{N_{\text{dof}}}$. $\Delta_{r\,n}$ is the (constant) value of $\Delta_{r}(a)$ between $a^{*}_{n-1}$ and $a^{*}_{n}$, while $a^{*}_{N_{\text{dof}}+1}=a_{\text{eq}}$ and the start of integration corresponds to $a^{*}_{0}=a_{0}(2H_{0}\Omega_{r}^{1/2}(\epsilon_{r}+1)t_{\text{pl}})^{1/2}/(G\mu)$ (corresponding to the start of significant GW emission from cosmic string loops). In this case, the total contribution of loops that decay during the radiation era may be rewritten as: